# Facility Location Games with Scaling Effects

### v1

@@ Abstract
We take the classic facility location problem and consider a variation, in which each agent’s individual cost function is equal to their distance from the facility multiplied by a scaling factor which is determined by the facility placement. In addition to the general class of continuous scaling functions, we also provide results for piecewise linear scaling functions which can effectively approximate or model the scaling of many real world scenarios. We focus on the objectives of total and maximum cost, describing the computation of the optimal solution. We then move to the approximate mechanism design setting, observing that the agents’ preferences may no longer be single-peaked. Consequently, we characterize the conditions on scaling functions which ensure that agents have single-peaked preferences. Under these conditions, we find results on the total and maximum cost approximation ratios achievable by strategyproof and anonymous mechanisms.
@@ 1 Introduction
  • When scaling functions meet the conditions for agents to have single-peaked preferences, we find the lower bound on the approximation ratios achieved by any strategic and anonymous mechanism, and propose such mechanisms with approximation ratios matching this lower bound. Table 1 summarizes some of our approximation ratio results.
General Preferences Single-Peaked Preferences
@@
Total Cost $r_{q}$ (Thm. 6.4) $e$ (Cor. 4) $(1+\frac{1}{k})^{k}$ (Cor. 5)
Max. Cost $2r_{q}$ (Thm. 6.6) $2e$ (Cor. 6) $2(1+\frac{1}{k})^{k}$ (Cor. 7)
@@ 2 Related Work
The facility location problem can be interpreted as a special case of preference aggregation when agents have symmetric and single-peaked preferences over a continuous range of outcomes, in which there has been decades of research in strategyproof aggregation rules. For general preferences, the famous Gibbard-Satthertwaite theorem implies that a surjective and strategyproof rule must be dictatorial [11, 22]. This impossibility result vanishes when we consider the restricted domain of single-peaked preferences, in which there have been numerous characterizations of strategyproof rules. Most notably, under this restricted domain, Moulin [17]’s seminal paper characterizes the space of strategyproof and anonymous rules as Phantom mechanisms which output the median of the agent locations and $n+1$ constant ‘Phantom’ values. This characterization was later extended to the domain of symmetric and single-peaked preferences by Border and Jordan [6] and Massó and Moreno De Barreda [15]. Other characterizations of strategyproof rules have been proposed for similar settings by [5, 18, 19, 13]. The paper most similar to our work considers the facility location model with entrance fees [14], in which agents incur, in addition to their distance from the facility, an additive cost which depends on the facility placement. In this model, the authors give strategyproof mechanisms and compute their approximation ratios with respect to the total and maximum cost objectives. The key difference with our model is that our agents experience a multiplicative scaling factor to their cost, whilst their agents experience an additional additive cost. Our work and [14] are inspired by the line of research on approximate mechanism design without money, initially proposed for the facility location problem by Procaccia and Tennenholtz [21]. In their paper, the authors compute the worst-case ratio between the performance of strategyproof mechanisms and welfare-optimal mechanisms. There have since been many papers applying this approach to variations of the facility location problem, such as when there are proportionally fair distance constraints [4], when facilities have capacity limits [3, 2], or when the facility is obnoxious (i.e., agents prefer to be far away from the facility) [8]. A recent survey of related work is given by Chan et al. [7].
@@ 3 Model
Denoting $\mathcal{Q}$ as the space of all scaling functions, facility location mechanism $f:\mathcal{Q}\times X^{n}\rightarrow X$ maps the agent location profile $\boldsymbol{x}=(x_{1},\dots,x_{n})$ to the location of a facility $y$. We define the cost incurred by agent $i$ as its distance from the facility multiplied by the scaling factor: $c_{i}(q,y):=q(y)|y-x_{i}|$. Finally, we define the total cost of an instance as $TC(q,y,\boldsymbol{x}):=\sum_{i}c_{i}(q,y)=q(y)\sum_{i}|y-x_{i}|$, and the maximum cost of an instance as $MC(q,y,\boldsymbol{x}):=\max_{i}c_{i}(q,y)=q(y)\max_{i}|y-x_{i}|$. Respectively, we denote the optimal facility location which minimizes the total cost (resp. maximum cost) as $y^{*}_{TC}$ (resp. $y^{*}_{MC}$). It is typically ideal for the mechanism output to be independent of the agents’ labelling, so we are concerned with mechanisms that satisfy anonymity.
@@ 4 Computing the Optimal Solution
Consider the location profile $\textbf{x}=(0,0.5,1)$ and the scaling function $q(y)=\begin{cases}1,&y\in[0,0.5],\\ 0.5,&y\in(0.5,1].\end{cases}$ The optimal facility locations for total cost and maximum cost are the right limit $y^{*}\rightarrow 0.5^{+}$, but not at $y^{*}=0.5$, so the optimal solutions are not well-defined.
@@ 4.1 Total Cost
Since $q$ is a linear function, it is of the form $q(y)=ay+b$, where $a\in\mathbb{R}$ and $b\in\mathbb{R}_{>0}$. If $a=0$, then $q(y)$ is constant, and thus the optimal facility location is on the median agent. [^3] If there are an even number of agents, we break the tie in favour of the leftmost agent in the optimal interval. Due to symmetry, it suffices to consider the case where the scaling function is increasing, i.e., $a>0$. From this assumption, we know that the optimal solution cannot lie to the right of $x_{n}$.
Since $q$ is increasing with $y$, its minimum value is on $y=0$. We denote $x_{0}:=0$, and we know the optimal solution $y^{*}$ is in the interval $[x_{0},x_{n}]$. Suppose the optimal solution $y^{*}$ is in the interval $[x_{n_{1}},x_{n_{1}+1}]$, where $n_{1}\in[0,n-1]$ denotes the number of agents on the left-hand side of $y^{*}$. Also, let $n_{2}=n-n_{1}$ denote the number of agents on the right-hand side of $y^{*}$. Recall that the total cost corresponding to $y^{*}$, $TC(q,y^{*},\textbf{x})$ is $q(y^{*})\sum_{i}|y^{*}-x_{i}|$. Considering the term $\sum_{i}|y^{*}-x_{i}|$, we have
@@
Now we know that the total cost function can either be a linear function or a quadratic function of $y^{*}$, depending on whether $a^{\prime}=0$, $a^{\prime}<0$, or $a^{\prime}>0$ (recalling that $a>0$). There are three cases: either the total cost function is a linear function ($a^{\prime}=0$), a quadratic function which opens downwards ($a^{\prime}<0$), or a quadratic function which opens upwards ($a^{\prime}>0$). For the first two cases, it is clear that the minimum value of the total cost function must lie on an endpoint of the interval $[x_{n_{1}},x_{n_{1}+1}]$. It remains to consider the last case ($a^{\prime}>0$), where the total cost function $q(y^{*})\sum_{i}|y^{*}-x_{i}|=aa^{\prime}y^{*}+(am+a^{\prime}b)y^{*}+bm$ opens upwards. Consider the total cost function on the larger domain of $\mathbb{R}$. Since $a^{\prime}=(n_{1}-n_{2})>0$, we know there exists $z_{1}\in\mathbb{R}$ such that $a^{\prime}z_{1}+m=0$. Also, since $a>0$, there exists $z_{2}\in\mathbb{R}$ such that $q(z_{2})=0$. As a result, we see that the total cost function is equal to zero at either one or two points on $\mathbb{R}$; we now show that both of these points lie to the left of $x_{n_{1}}$. For the term $a^{\prime}y^{*}+m$, we know that $a^{\prime}z_{1}+m=0$, and thus $z_{1}=-\frac{m}{a^{\prime}}=\frac{\sum_{i=n_{1}+1}^{n}x_{i}-\sum_{i=1}^{n_{1}}x_{i}}{n_{2}-n_{1}}.$ Consider the numerator $\sum_{i=n_{1}+1}^{n}x_{i}-\sum_{i=1}^{n_{1}}x_{i}$. We know there are $n_{2}$ agents between $(x_{n_{1}},x_{n}]$, implying $\sum_{i=n_{1}+1}^{n}x_{i}>n_{2}\cdot x_{n_{1}}$. Similarly, we know there are $n_{1}$ agents between $[0,x_{n_{1}}]$, implying $\sum_{i=1}^{n_{1}}x_{i}\leq n_{1}\cdot x_{n_{1}}$. Combining these inequalities gives us
@@
Since $z_{1},z_{2}<x_{n_{1}}$, the total cost function must be strictly increasing in $[x_{n_{1}},x_{n_{1}+1}]$, so its minimum value must lie on $x_{n_{1}}$, which is either an agent location or the minimum value of $q$.
@@
First, consider each line segment on $q$, which is represented by a linear function $q_{j}$ on the sub-domain $[c_{j},d_{j}]$, where $c_{j},d_{j}\in X$. Let $y_{j}^{*}:=\arg\min_{y_{j}\in[c_{j},d_{j}]}TC(q,y_{j},\boldsymbol{x})$ denote the optimal facility placement on $[c_{j},d_{j}]$, and $e_{j}:=\arg\min_{y\in\{c_{j},d_{j}\}}q(y)$ be the minimum point of $q$ on $[c_{j},d_{j}]$. By Proposition 2, we must have $y_{j}^{*}\in\{e_{j},x_{1},\dots,x_{n}\}$. Therefore, considering all of the $q_{j}$ functions, the optimal solution $y^{*}$ must lie at one of the agents’ positions or at a local minimum of $q_{j}$.
@@
Suppose the scaling function is $q(y)=1-\frac{4}{5}y$ and the agent location profile is $x=(0.3,0.4,0.7)$. The optimal facility placement is $y=0.4$, which leads to a cost of $0.204$ for agent $x_{3}$. However if agent $x_{3}$ misreports $x_{3}^{\prime}=0.755$, then the facility placement is instead $y^{\prime}=1$, which leads to a cost of $0.06$. Hence the optimal facility placement is not strategyproof.
@@ 4.2 Maximum Cost
We now know that if $a>0$, the maximum cost is $MC(q,y,x)=c_{n}(q,y)=(ay+b)(x_{n}-y)$, which is a quadratic function of $y$ that opens downwards. Therefore the maximum cost is minimized when $y=0$ or $y=\frac{x_{1}+x_{n}}{2}$. By symmetry, if $a<0$, agent $1$ incurs the maximum cost of $MC(q,y,x)=c_{1}(q,y)=(ay+b)(y-x_{1})$, which is minimized when $y=1$ or $y=\frac{x_{1}+x_{n}}{2}$. We therefore have $y^{*}_{MC}\in\{0,\frac{x_{1}+x_{n}}{2},1\}$.
@@
$y^{*}_{MC}\in\left\{\operatorname*{arg\,min}_{y\in[0,\frac{x_{1}+x_{n}}{2}]}c_{n}(q,y),\operatorname*{arg\,min}_{y\in[\frac{x_{1}+x_{n}}{2},1]}c_{1}(q,y),\frac{x_{1}+x_{n}}{2}\right\}.$
@@
If $y\in[0,\frac{x_{1}+x_{n}}{2}]$, then agent $n$ receives the highest cost in the instance, similar to the proof of Proposition 4. A symmetric argument shows that if $y\in[\frac{x_{1}+x_{n}}{2},1]$, then agent $1$ receives the highest cost in the instance. Therefore $y^{*}_{MC}$ is either on the minimum value of $c_{n}(q,y)$ on $y\in[0,\frac{x_{1}+x_{n}}{2}]$, the minimum value of $c_{1}(q,y)$ on $y\in[\frac{x_{1}+x_{n}}{2},1]$, or on $\frac{x_{1}+x_{n}}{2}$.
@@ 5 Achieving Single-Peaked Preferences
When agents have single-peaked preferences along a real line domain, Moulin [17] characterized the set of anonymous and strategyproof mechanisms as phantom mechanisms, which place the facility at the median of the $n$ agent locations and $n+1$ constant ‘phantom’ points.
@@
The median mechanism places the facility at $y=0.7$, and it leads to a cost of $0.6$ for agent $x_{3}$. However if the agent at $x_{3}$ misreports to $x_{3}^{\prime}=0.5$, then the median facility placement will be $0.5$, which leads to a lower cost of $0.4$ for the agent at $x_{3}$. Hence the median mechanism is not strategyproof.
@@
Therefore, a phantom mechanism which does not place all phantoms at the same position is not strategyproof.
@@
The constant mechanism (which places the facility at some constant $c\in X$) is the only anonymous and strategyproof mechanism in our setting.
@@
Consider the scaling function $q(y)=y+0.01$ and $x_{1}=0.5$. The derivative of agent $1$’s utility is $u_{1}^{\prime}(q,y)=2y-0.49$ when $y<0.5$, and $u_{1}^{\prime}(q,y)=-(2y-0.49)$ when $y>0.5$. For $y\in[0,0.245]$, $u_{1}^{\prime}(q,y)\leq 0$. For $y\in[0.245,0.5]$, $u^{\prime}\geq 0$. Lastly, for $y\in[0.5,1]$, $u_{1}^{\prime}(q,y)<0$. Therefore $u_{1}(q,y)$ is not single-peaked.
@@
From Theorem 5.3, we know that if the $q$ function is continuous but not differentiable at some set of points $D$, the utility function is single-peaked if and only if $q(y)$ satisfies $q(y)-|q^{\prime}(y)|\geq 0$ for all $y\in[0,1]\backslash D$. When $a_{j}<0$, the utility function is single-peaked if and only if $a_{j}y+b_{j}+a_{j}\geq 0\Rightarrow b_{j}\geq-2a_{j}y$, and when $a_{j}>0$, the utility function is single-peaked if and only if $a_{j}y+b_{j}-a_{j}\geq 0\Rightarrow b_{j}\geq a_{j}(y-1)$. Since $y\in[0,1]$, these conditions become $b_{j}\geq-2a_{j}$ and $b_{j}\geq a_{j}$, respectively.

### v2

@@ Abstract
We take the classic facility location problem and consider a variation, in which each agent’s individual cost function is equal to their distance from the facility multiplied by a scaling factor which is determined by the facility placement. In addition to the general class of continuous scaling functions, we also provide results for piecewise linear scaling functions which can effectively approximate or model the scaling of many real world scenarios. We focus on the objectives of total and maximum cost, describing the computation of the optimal solution. We then move to the approximate mechanism design setting, observing that the agents’ preferences may no longer be single-peaked. Consequently, we characterize the conditions on scaling functions which ensure that agents have single-peaked preferences. Under these conditions, we find a characterization of continuous, strategyproof, and anonymous mechanisms, and compute the total and maximum cost approximation ratios achievable by these mechanisms.
@@ 1 Introduction
  • When scaling functions meet the conditions for agents to have single-peaked preferences, we characterize the space of continuous mechanisms satisfying strategyproofness and anonymity as phantom mechanisms with scaling.
  • We find the lower bound on the approximation ratios achieved by phantom mechanisms with scaling, and propose such mechanisms with approximation ratios matching this lower bound. Table 1 summarizes some of our approximation ratio results.
General Pref. Single-Peaked Pref.
@@
Total Cost $r_{q}$ (Thm. 7.3) $e$ (Cor. 4) $(1+\frac{1}{k})^{k}$ (Cor. 5)
Max. Cost $2r_{q}$ (Thm. 7.5) $2e$ (Cor. 6) $2(1+\frac{1}{k})^{k}$ (Cor. 7)
@@ 2 Related Work
The facility location problem can be interpreted as a special case of preference aggregation when agents have symmetric and single-peaked preferences over a continuous range of outcomes, in which there has been decades of research in strategyproof aggregation rules. For general preferences, the famous Gibbard-Satthertwaite theorem implies that a surjective and strategyproof rule must be dictatorial [12, 23]. This impossibility result vanishes when we consider the restricted domain of single-peaked preferences, in which there have been numerous characterizations of strategyproof rules. Most notably, under this restricted domain, Moulin [18]’s seminal paper characterizes the space of strategyproof and anonymous rules as Phantom mechanisms which output the median of the agent locations and $n+1$ constant ‘Phantom’ values. This characterization was later extended to the domain of symmetric and single-peaked preferences by Border and Jordan [6] and Massó and Moreno De Barreda [16]. Other characterizations of strategyproof rules have been proposed for similar settings by [5, 19, 20, 14].
The paper most similar to our work considers the facility location model with entrance fees [15], in which agents incur, in addition to their distance from the facility, an additive cost which depends on the facility placement. In this model, the authors give strategyproof mechanisms and compute their approximation ratios with respect to the total and maximum cost objectives. The key difference with our model is that our agents experience a multiplicative scaling factor to their cost, whilst their agents experience an additional additive cost. Our work and [15] are inspired by the line of research on approximate mechanism design without money, initially proposed for the facility location problem by Procaccia and Tennenholtz [22]. In their paper, the authors compute the worst-case ratio between the performance of strategyproof mechanisms and welfare-optimal mechanisms. There have since been many papers applying this approach to variations of the facility location problem, such as when there are proportionally fair distance constraints [4], when facilities have capacity limits [3, 2], or when the facility is obnoxious (i.e., agents prefer to be far away from the facility) [9]. A recent survey of related work is given by Chan et al. [8].
@@ 3 Model
Denoting $\mathcal{Q}$ as the space of all scaling functions, a continuous [^3] For mechanisms, we define continuity with respect to the agents’ locations. Formally, we say a mechanism is continuous if $\forall\mathbf{x}\in X^{n}:\forall\epsilon>0:\exists\delta>0:\forall\mathbf{x}^{\prime}\in X^{n},\forall q\in\mathcal{Q}:\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert_{1}<\delta\implies\lVert f(q,\mathbf{x})-f(q,\mathbf{x}^{\prime})\rVert_{1}<\epsilon$. facility location mechanism $f:\mathcal{Q}\times X^{n}\rightarrow X$ maps the agent location profile $\boldsymbol{x}=(x_{1},\dots,x_{n})$ to the location of a facility $y$. We define the cost incurred by agent $i$ as its distance from the facility multiplied by the scaling factor: $c_{i}(q,y):=q(y)|y-x_{i}|$. Finally, we define the total cost of an instance as $TC(q,y,\boldsymbol{x}):=\sum_{i}c_{i}(q,y)=q(y)\sum_{i}|y-x_{i}|$, and the maximum cost of an instance as $MC(q,y,\boldsymbol{x}):=\max_{i}c_{i}(q,y)=q(y)\max_{i}|y-x_{i}|$. Respectively, we denote the optimal facility location which minimizes the total cost (resp. maximum cost) as $y^{*}_{TC}$ (resp. $y^{*}_{MC}$).
It is typically ideal for the mechanism output to be independent of the agents’ labelling, so we are concerned with mechanisms that satisfy anonymity.
@@ 4 Computing the Optimal Solution
Consider the location profile $\textbf{x}=(0,0.5,1)$ and the scaling function $q(y)=\begin{cases}1,&y\in[0,0.5],\\ 0.5,&y\in(0.5,1].\end{cases}$ The optimal facility locations for total cost and maximum cost are the right limit $y^{*}\rightarrow 0.5^{+}$, but not at $y^{*}=0.5$, so the optimal solutions are not well-defined. ∎
@@ 4.1 Total Cost
Since $q$ is a linear function, it is of the form $q(y)=ay+b$, where $a\in\mathbb{R}$ and $b\in\mathbb{R}_{>0}$. If $a=0$, then $q(y)$ is constant, and thus the optimal facility location is on the median agent. [^4] If there are an even number of agents, we break the tie in favour of the leftmost agent in the optimal interval.
Due to symmetry, it suffices to consider the case where the scaling function is increasing, i.e., $a>0$. From this assumption, we know that the optimal solution cannot lie to the right of $x_{n}$.
Since $q$ is increasing with $y$, its minimum value is on $y=0$. We denote $x_{0}:=0$, and we know the optimal solution $y^{*}$ is in the interval $[x_{0},x_{n}]$. Suppose the optimal solution $y^{*}$ is in the interval $[x_{n_{1}},x_{n_{1}+1}]$, where $n_{1}\in[0,n-1]$ denotes the number of agents on the left-hand side of $y^{*}$. Also, let $n_{2}=n-n_{1}$ denote the number of agents on the right-hand side of $y^{*}$.
Recall that the total cost corresponding to $y^{*}$, $TC(q,y^{*},\textbf{x})$ is $q(y^{*})\sum_{i}|y^{*}-x_{i}|$. Considering the term $\sum_{i}|y^{*}-x_{i}|$, we have
@@
Now we know that the total cost function can either be a linear function or a quadratic function of $y^{*}$, depending on whether $a^{\prime}=0$, $a^{\prime}<0$, or $a^{\prime}>0$ (recalling that $a>0$). There are three cases: either the total cost function is a linear function ($a^{\prime}=0$), a quadratic function which opens downwards ($a^{\prime}<0$), or a quadratic function which opens upwards ($a^{\prime}>0$). For the first two cases, it is clear that the minimum value of the total cost function must lie on an endpoint of the interval $[x_{n_{1}},x_{n_{1}+1}]$. It remains to consider the last case ($a^{\prime}>0$), where the total cost function $q(y^{*})\sum_{i}|y^{*}-x_{i}|=aa^{\prime}y^{*}+(am+a^{\prime}b)y^{*}+bm$ opens upwards. Consider the total cost function on the larger domain of $\mathbb{R}$. Since $a^{\prime}=(n_{1}-n_{2})>0$, we know there exists $z_{1}\in\mathbb{R}$ such that $a^{\prime}z_{1}+m=0$. Also, since $a>0$, there exists $z_{2}\in\mathbb{R}$ such that $q(z_{2})=0$. As a result, we see that the total cost function is equal to zero at either one or two points on $\mathbb{R}$; we now show that both of these points lie to the left of $x_{n_{1}}$.
For the term $a^{\prime}y^{*}+m$, we know that $a^{\prime}z_{1}+m=0$, and thus $z_{1}=-\frac{m}{a^{\prime}}=\frac{\sum_{i=n_{1}+1}^{n}x_{i}-\sum_{i=1}^{n_{1}}x_{i}}{n_{2}-n_{1}}.$ Consider the numerator $\sum_{i=n_{1}+1}^{n}x_{i}-\sum_{i=1}^{n_{1}}x_{i}$. We know there are $n_{2}$ agents between $(x_{n_{1}},x_{n}]$, implying $\sum_{i=n_{1}+1}^{n}x_{i}>n_{2}\cdot x_{n_{1}}$. Similarly, we know there are $n_{1}$ agents between $[0,x_{n_{1}}]$, implying $\sum_{i=1}^{n_{1}}x_{i}\leq n_{1}\cdot x_{n_{1}}$. Combining these inequalities gives us
@@
Since $z_{1},z_{2}<x_{n_{1}}$, the total cost function must be strictly increasing in $[x_{n_{1}},x_{n_{1}+1}]$, so its minimum value must lie on $x_{n_{1}}$, which is either an agent location or the minimum value of $q$. ∎
@@
First, consider each line segment on $q$, which is represented by a linear function $q_{j}$ on the sub-domain $[c_{j},d_{j}]$, where $c_{j},d_{j}\in X$. Let $y_{j}^{*}:=\arg\min_{y_{j}\in[c_{j},d_{j}]}TC(q,y_{j},\boldsymbol{x})$ denote the optimal facility placement on $[c_{j},d_{j}]$, and $e_{j}:=\arg\min_{y\in\{c_{j},d_{j}\}}q(y)$ be the minimum point of $q$ on $[c_{j},d_{j}]$. By Proposition 2, we must have $y_{j}^{*}\in\{e_{j},x_{1},\dots,x_{n}\}$. Therefore, considering all of the $q_{j}$ functions, the optimal solution $y^{*}$ must lie at one of the agents’ positions or at a local minimum of $q_{j}$. ∎
@@
Suppose the scaling function is $q(y)=1-\frac{4}{5}y$ and the agent location profile is $x=(0.3,0.4,0.7)$. The optimal facility placement is $y=0.4$, which leads to a cost of $0.204$ for agent $x_{3}$. However, if agent $x_{3}$ misreports $x_{3}^{\prime}=0.755$, then the facility placement is instead $y^{\prime}=1$, which leads to a cost of $0.06$. Hence the optimal facility placement is not strategyproof. ∎
@@ 4.2 Maximum Cost
We now know that if $a>0$, the maximum cost is $MC(q,y,x)=c_{n}(q,y)=(ay+b)(x_{n}-y)$, which is a quadratic function of $y$ that opens downwards. Therefore the maximum cost is minimized when $y=0$ or $y=\frac{x_{1}+x_{n}}{2}$. By symmetry, if $a<0$, agent $1$ incurs the maximum cost of $MC(q,y,x)=c_{1}(q,y)=(ay+b)(y-x_{1})$, which is minimized when $y=1$ or $y=\frac{x_{1}+x_{n}}{2}$. We therefore have $y^{*}_{MC}\in\{0,\frac{x_{1}+x_{n}}{2},1\}$. ∎
@@
$y^{*}_{MC}\in\left\{\operatorname*{arg\,min}_{y\in[0,\frac{x_{1}+x_{n}}{2}]}c_{n}(q,y),\operatorname*{arg\,min}_{y\in[\frac{x_{1}+x_{n}}{2},1]}c_{1}(q,y)\right\}.$
@@
If $y\in[0,\frac{x_{1}+x_{n}}{2}]$, then agent $n$ receives the highest cost in the instance, similar to the proof of Proposition 4. A symmetric argument shows that if $y\in[\frac{x_{1}+x_{n}}{2},1]$, then agent $1$ receives the highest cost in the instance. Therefore $y^{*}_{MC}$ is either on the minimum value of $c_{n}(q,y)$ on $y\in[0,\frac{x_{1}+x_{n}}{2}]$, or the minimum value of $c_{1}(q,y)$ on $y\in[\frac{x_{1}+x_{n}}{2},1]$. ∎
@@ 5 Achieving Single-Peaked Preferences
When agents have single-peaked preferences along a compact domain, Moulin [18] characterized the set of anonymous and strategyproof mechanisms as phantom mechanisms, which place the facility at the median of the $n$ agent locations and $n+1$ constant ‘phantom’ points.
@@
The median mechanism places the facility at $y=0.7$, and it leads to a cost of $0.6$ for agent $x_{3}$. However if the agent at $x_{3}$ misreports to $x_{3}^{\prime}=0.5$, then the median facility placement will be $0.5$, which leads to a lower cost of $0.4$ for the agent at $x_{3}$. Hence the median mechanism is not strategyproof. ∎
@@
Therefore, a phantom mechanism which does not place all phantoms at the same position is not strategyproof. ∎
@@
The constant mechanism (which places the facility at some constant $c\in X$) is the only anonymous and strategyproof phantom mechanism in our setting.
@@
Consider the scaling function $q(y)=y+0.01$ and $x_{1}=0.5$. The derivative of agent $1$’s utility is $u_{1}^{\prime}(q,y)=2y-0.49$ when $y<0.5$, and $u_{1}^{\prime}(q,y)=-(2y-0.49)$ when $y>0.5$. For $y\in[0,0.245]$, $u_{1}^{\prime}(q,y)\leq 0$. For $y\in[0.245,0.5]$, $u^{\prime}\geq 0$. Lastly, for $y\in[0.5,1]$, $u_{1}^{\prime}(q,y)<0$. Therefore $u_{1}(q,y)$ is not single-peaked. ∎
@@
From Theorem 5.3, we know that if the $q$ function is continuous but not differentiable at some set of points $D$, the utility function is single-peaked if and only if $q(y)$ satisfies $q(y)-|q^{\prime}(y)|\geq 0$ for all $y\in[0,1]\backslash D$. When $a_{j}<0$, the utility function is single-peaked if and only if $a_{j}y+b_{j}+a_{j}\geq 0\Rightarrow b_{j}\geq-2a_{j}y$, and when $a_{j}>0$, the utility function is single-peaked if and only if $a_{j}y+b_{j}-a_{j}\geq 0\Rightarrow b_{j}\geq a_{j}(y-1)$. Since $y\in[0,1]$, these conditions become $b_{j}\geq-2a_{j}$ and $b_{j}\geq a_{j}$, respectively. ∎
6 Characterization of Strategyproof and Anonymous Mechanisms
As we have characterized the conditions for the scaling function to guarantee single-peaked agent preferences, it may seem immediate to apply the results by Moulin [18] to characterize strategyproof and anonymous mechanisms under these conditions as phantom mechanisms. However, recall that the mechanism takes the scaling function as an additional input, meaning that the constant value $p_{1},\dots,p_{n+1}$ may be dependent on the scaling function. Furthermore, the domain of single-peaked preferences induced by a scaling function meeting the key condition of Theorem 5.3 may be a strict subset of the domain of all arbitrary single-peaked preferences. As a result, the characterization by Moulin [18] does not immediately hold in our setting.
Nevertheless, we are still able to obtain a similar characterization of strategyproof and anonymous mechanisms in our setting when the scaling function guarantees single-peaked agent preferences, but we additionally require that the mechanism is continuous. We first define an adaptation of the phantom mechanism, in which each ‘constant’ value admits the scaling function as input.
Definition 4 (Phantom Mechanism with Scaling)
Given an agent location profile $\mathbf{x}$, a scaling function $q$, and $n+1$ ‘phantom values’ $\{p_{i}(q)\}_{i\in[n+1]}$ defined by $p_{i}:\mathcal{Q}\rightarrow[0,1]$, a phantom mechanism with scaling places the facility at
$\operatorname{med}\{x_{1},\cdots,x_{n},p_{1}(q),\cdots,p_{n+1}(q)\}.$
We now present our characterization, which leverages a key result by Brandt et al. [7].
Theorem 6.1
When agents are guaranteed to have single-peaked preferences, a continuous mechanism is strategyproof and anonymous if and only if it is a phantom mechanism with scaling.
Proof
[…]
()($\implies$)( ⟹ ) Let $\mathcal{U}^{SP}$ denote the set of all single-peaked preferences. Theorem 1 of [7] states that for an arbitrary $\mathcal{U}\subseteq\mathcal{U}^{SP}$, a continuous mechanism $f$ is anonymous and strategyproof if and only if there exists $p_{1},\dots,p_{n+1}$ in $[0,1]$ such that the facility is placed at $\text{med}\{x_{1},\dots,x_{n},p_{1},\dots,p_{n+1}\}$. Note that $p_{1},\dots,p_{n+1}$ must be independent of the agents’ locations, but can depend on the preference domain $\mathcal{U}$. In our model, the set of single-peaked preferences $\mathcal{U}$ is induced by the scaling function, and hence $p_{1},\dots,p_{n+1}$ can be dependent on the scaling function. It immediately follows that when agents have single-peaked preferences, a continuous mechanism which is strategyproof and anonymous must be a phantom mechanism with scaling. ∎